OpenStudy (anonymous):

HELP!!!!! Word problem using 2 variables & 2 equations. "As of 2007, the NY Yankees & the Cincinnati Reds together have won 31 World Series. The Yankees had won 5.2 times as many as the Reds. How many World Series had each team won? X = # of time Yankees won Y = # of time Reds won Eq1: x + y = 31(I'm confused about Eq#2) Eq2:

OpenStudy (mathsolver):

5.2Y=X

OpenStudy (cwrw238):

x = 5.2y

OpenStudy (mathsolver):

I like to see we're on the same tracks @cwrw238

OpenStudy (cwrw238):

yep

OpenStudy (anonymous):

yes. Let me try to solve now. I was confused as to how to set up eq 2

OpenStudy (anonymous):

I got it - THANKS. The Yankees won 26 world series and the Reds 5. Here is how I solved. Eq #1: x+y=31 Eq #2: x=5.2y y+(5.2y)=31 6.2y=31 divide both sides by 6.2 and y=5 then to solve for x we do the Eq #1. x+5=31 subtract 5 to both sides and x=26

OpenStudy (mathsolver):

well done!

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Thank you both - Unknown how to give out two medals?

OpenStudy (mathsolver):

don't worry about it. I just gave @cwrw238 a medal on your behalf

OpenStudy (anonymous):

thank you!